OpenStudy (anonymous):

Which algebraic expression is equivalent to the expression below? (5x + 12) + 6x 23x (5x + 6x) + 12 (5x - 6x) + 12 (5x + 12x) + 6

OpenStudy (anonymous):

the second , because with adding operation u can use commutative propety (you don't even need parenthesis); the result however is 11x+12

OpenStudy (unklerhaukus):

(A+ B) + C = A+ (B + C) (by the associative property of addition) = A+ (C + B) (by the commutative property of addition) =(A+ C) + B (by the associative property of addition)